Upton Umbrellas has a cost of equity of 12.4 percent, the YTM on the company's bonds is 5.9 percent, and the tax rate is 40 percent. The company's bonds sell for 103.8 percent of par. The debt has a book value of $432,000 and total assets have a book value of $960,000. If the market-to-book ratio is 2.98 times, what is the company's WACC?
Answer #1

Total book value of equity=Total assets book value-Total debt book value

=(960000-432000)= 528,000

market-to-book ratio=market value/book value

Hence market value of equity=(2.98*528000)= 1573440

After tax cost of debt= YTM (1- tax rate )= 5.9% (1-0.4)

= 3.54%

Market value Cost
Equity 1573440 12.40%
Debt (432000*103.80%)=448416 3.54%
Total 2021856

WACC= Equity / Total market value * cost of Equity + Debt /total market value * cost of Debt

=1573440/2021856*12.40% +448416/2021856*3.54%

= 9.6499% + 0.78512% = 10.44%